OpenStudy (anonymous):

How many solutions does the following system of equations have x = −2 and x2 + y2 = 4

OpenStudy (anonymous):

y^2 + 4 = 4 y^2 = 0 y=0 If y was a number other than zero, then it would be +/- that number (two solutions

OpenStudy (anonymous):

So say y^2 = 9, then y =3 and y=-3

sam (.sam.):

x2 + y2 = 4 when x=-2, \[(-2)^2 + y^2=4\] \[4+y^2=4\] \[y^2=2\] \[y=\pm \sqrt{2}\] ---------------------- when \[y=+\sqrt{2}\] \[x^2 +2=4\] \[x^2 =2\] \[x=\pm \sqrt{2}\] --------------------- when\[ y=-\sqrt{2}\] \[x=\pm \sqrt{2}\] =============== 4 solutions

OpenStudy (anonymous):

@.Sam. maybe I'm wrong, but: (-2)^2+y^2 = 4 4 + y^2 = 4 -4 + 4 + y^2 = 4 -4 y^2 = 0 does not = 2